OpenStudy (anonymous):

how can I factor (3n-2)(4n+1)=0 I already know that I have to find a,b,c but theres no sign in between those two parentheses, what do I do?

OpenStudy (anonymous):

It is factored. Are you trying to solve for n?

OpenStudy (anonymous):

maybe he is trying to multiply them to put it back into the form: \[an^2+bn+c\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes to which one?

OpenStudy (anonymous):

Im trying to solve it in quadratic formula way

OpenStudy (anonymous):

but theres no addition or sub. in between (3n-2)(4n+1)

OpenStudy (anonymous):

If you are looking to use the quadratic formula then you would have to multiply the two binomials first. It would be faster to solve for n without doing so.

OpenStudy (anonymous):

i think you are making the problem a little harder than it has to be. Unless the teacher specifically asks for you to use the quadratic formula, if it looks like the equation above i wouldnt do it that way.

OpenStudy (anonymous):

hmmm ok so how exactly would you answer this question because I got 2/3, 1/4 but how they got it I dont know

OpenStudy (anonymous):

You can solve for n by taking each binomial and setting it equal to zero. Then solving for n. For example: 3n-2 = 0 4n +1 =0 Solve those for n and you will have the possible solutions.

OpenStudy (anonymous):

k thanks and how about n2= -18-9n

OpenStudy (anonymous):

and -2v2-v+12=-3v2+6v step by step please

OpenStudy (anonymous):

if that is \[n ^{2} =-18-9n\] then you would need to have all the variables on one side and you could use the quadratic formula.

OpenStudy (anonymous):

on one side? like -18-9n=n2?

OpenStudy (anonymous):

all on one side like \[n ^{2}+9n+18=0\]

OpenStudy (anonymous):

does it automatically become a plus ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

it became positive because we added it to both sides of the equation.